OpenStudy (wwe123):

if the mth,nth and pth terms of an AP and GP are equal and are x,y,z respectively,prove that

OpenStudy (wwe123):

prove that \[x ^{y-z}.y ^{z-x}.z ^{x-y}=1\]

OpenStudy (experimentx):

define each of them first x = a + (m-1)d <-- for ap, = a r^(m-1) <--- for gp

OpenStudy (experimentx):

woops! x = a + (m-1)d <-- for ap, = b r^(m-1) <--- for gp

OpenStudy (wwe123):

plz explain more

OpenStudy (experimentx):

i don't know at the moment, but i would make full use of info given towards the problem.

OpenStudy (wwe123):

plz solve the question

OpenStudy (experimentx):

okay, write the denominators are the Geometric terms and the powers are the Arithemetic terms!! the power cancel up to give zero ... hence you get RHS

OpenStudy (experimentx):

sorry ... *write the bases as Geometric terms and exponent as Arithmetic terms

OpenStudy (wwe123):

just show it...

OpenStudy (experimentx):

 do the rest for others and multiply it.